darkknight:

Math question You can prove 1 = 2

darkknight:

let a variable a = b multiply both sides by a \[a^2=ab\] square subtract by b^2 on both sides \[a^2-b^2=ab - b^2\] the left side of the eqn is a difference of squares \[(a+b)(a-b) = b(a-b)\] We can divide both sides by a-b \[a+b = b\] Remember a = b, so \[2b=b\] divide both sides by b \[2=1\]!?!?!?!

darkknight:

Math question, I would like to hear if you think this is right, or if wrong how come. ¯\_(ツ)_/¯ @mxddi3 @justus @ramen

ramen:

I can tell you were bored, but lemme see

AZ:

This is a classic example, however, the only problem is that since we assume a = b, that means a-b = 0 We can't divide both sides by 0

darkknight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ This is a classic example, however, the only problem is that since we assume a = b, that means a-b = 0 We can't divide both sides by 0 \(\color{#0cbb34}{\text{End of Quote}}\) xD yea u kinda got it @ljtheballer16 remember when we divided both sides by a-b? so we know a = b so basically we divided both sides by either a-a (if we substitute a for b) or b-b (by substituting b for a), both of these equal 0. What I did between step 3 and 4 was divide both sides by a-b, so this is actually not valid and is a trick, Nice job @AZ for figuring it out!

LjTheBaller16:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight \(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ This is a classic example, however, the only problem is that since we assume a = b, that means a-b = 0 We can't divide both sides by 0 \(\color{#0cbb34}{\text{End of Quote}}\) xD yea u kinda got it @ljtheballer16 remember when we divided both sides by a-b? so we know a = b so basically we divided both sides by either a-a (if we substitute a for b) or b-b (by substituting b for a), both of these equal 0. What I did between step 3 and 4 was divide both sides by a-b, so this is actually not valid and is a trick, Nice job @AZ for figuring it out! \(\color{#0cbb34}{\text{End of Quote}}\) ohhh i gotchu